# RiemannLoRA: A Unified Riemannian Framework for Ambiguity-Free LoRA Optimization

### v1

@@ 5 Experiments
In this section, we also employ the term LOI (Locally Optimal Initialization), referring to the proposed initialization scheme described in Section 4.2. This is done to distinguish between using RiemannLoRA as an optimizer with zero initialization (as in basic LoRA) and RiemannLoRA-LOI, which is the proposed combination of initialization and optimization. We conduct a series of LLM fine-tuning experiments for different tasks. All of the experiments were computed on NVIDIA V100-SXM2-32Gb GPU and NVIDIA A100-SXM-80Gb GPU. We ran all the experiments within $\sim 2000$ GPU hours.
@@ 5.1 Commonsense reasoning fine-tuning
The results were obtained for the benchmark (Clark et al. (2019, BoolQ), Bisk et al. (2020, PIQA), Sap et al. (2019, SIQA), Zellers et al. (2019, hellaswag), Sakaguchi et al. (2021, winogrande),Clark et al. (2018, ARC), Mihaylov et al. (2018, OBQA)) common reasoning. The structure of the dataset is described in Appendix B. In the following experiments we conduct a fine-tuning procedures for multilayer perceptron (MLP) and attention layers of Llama 3.2 1b model (Dubey et al. (2024)).
@@
5.1.1 Fine-tuning optimizers comparison
For every fine-tuning method we preselected a suitable optimization step sizes. For RiemannLoRA Adam (see the Algorithm 3 with simulate_Adam flag enabled) we also selected the hyperparameter $\beta$, $\gamma$. The table 1 contains the accuracy of the trained via SGD without Heavy-Ball momentum model’s responses on the test dataset. In the frame of SGD-like optimizers the exploitation of the proposed Riemannian fine-tuning framework demonstrates a performance increase with and without local optimal initialization: the RiemannLoRA with local optimal initialization outperforms LoRA, DoRA, RSLoRA. The standard derivation of the presented method is the least among all other methods.
Table 2 contains results for Adam-like optimizers. The proposed method demonstrates the best (together with LoRA-LOI) mean metric value. It is also notable that our method produces least variance compared with all other approaches. Note also, that our Riemannian version is for a simplified Adam optimizer. A more advanced version, better aligned with Adam is a subject for future work.
Table 1: The average accuracy (in %) among $8$ tasks of fine-tuned Llama 3.2-1b using different SGD variants, tested on Commonsense Reasoning benchmark. The <<R->> prefix stands for Riemannian, the postfix <<-LOI>> stands for locally optimal initialization. LoRA rank is set to 16. In all experiments excpet for <<-LOI>> the $A$ factor in $AB^{\top}$ has orthonormal columns and $B$ is zero.
@@ 5.1.1 Fine-tuning optimizers comparison
Raw $40.1$ $55.4$ $50.3$ $25.8$ $50.0$ $61.9$ $41.8$ $42.8$ $46.0$
RSLoRA $\mathbf{64.5}_{\pm 0.2}$ $77.2_{\pm 0.3}$ $68.1_{\pm 0.8}$ $86.3_{\pm 0.1}$ $58.2_{\pm 0.5}$ $76.5_{\pm 1.5}$ $58.6_{\pm 2.9}$ $70.0_{\pm 1.9}$ $70.0_{\pm 0.8}$
RiemannLoRA $65.1_{\pm 0.4}$ $77.7_{\pm 0.3}$ $69.1_{\pm 1.5}$ $85.8_{\pm 0.2}$ $58.5_{\pm 0.4}$ $74.4_{\pm 1.9}$ $56.9_{\pm 1.3}$ $69.3_{\pm 2.5}$ $69.6_{\pm 0.9}$
RiemannLoRA-LOI $65.2_{\pm 0.4}$ $\mathbf{79.4}_{\pm 0.4}$ $\mathbf{75.6}_{\pm 0.4}$ $\mathbf{87.3}_{\pm 0.1}$ $\mathbf{62.4}_{\pm 0.4}$ $\mathbf{79.9}_{\pm 0.8}$ $\mathbf{63.6}_{\pm 1.9}$ $\mathbf{73.8}_{\pm 0.5}$ $\mathbf{73.4}_{\pm 0.3}$
Figure 1: The train loss of fine-tuning Llama 3.2-1b vs optimization step using different SGD variants on Commonsense Reasoning benchmark. The <<R->> prefix stands for Riemannian, the postfix <<-LOI>> stands for locally optimal initialization. LoRA rank is set to 16. In all experiments excpet for <<-LOI>> the $A$ factor in $AB^{\top}$ has orthonormal columns and $B$ is zero.
Table 2: The average accuracy (in %) among $8$ tasks of fine-tuned Llama 3.2-1b using different Adam variants, tested on Commonsense Reasoning benchmark. The <<R->> prefix stands for Riemannian, the postfix <<-LOI>> stands for locally optimal initialization. LoRA rank is set to 16.
Task BoolQ PIQA SIQA hella- wino- ARC-E ARC-C OBQA All
Initialization swag grande
Raw $40.1$ $55.4$ $50.3$ $25.8$ $50.0$ $61.9$ $41.8$ $42.8$ $46.0$
LoRA $65.4_{\pm 1.7}$ $79.5_{\pm 1.1}$ $74.9_{\pm 0.3}$ $87.2_{\pm 0.5}$ $62.4_{\pm 1.8}$ $\mathbf{79.3}_{\pm 0.6}$ $62.3_{\pm 1.2}$ $75.5_{\pm 0.5}$ $73.3_{\pm 0.7}$
LoRA-LOI $65.6_{\pm 1.4}$ $\mathbf{81.3}_{\pm 0.9}$ $75.7_{\pm 0.6}$ $87.7_{\pm 0.6}$ $\mathbf{65.7}_{\pm 2.4}$ $77.9_{\pm 1.0}$ $65.0_{\pm 1.7}$ $75.2_{\pm 0.9}$ $\mathbf{74.3}_{\pm 0.9}$
LoRA-QR $\mathbf{66.0}_{\pm 1.8}$ $79.9_{\pm 0.8}$ $74.6_{\pm 0.6}$ $87.5_{\pm 0.6}$ $63.5_{\pm 2.6}$ $79.0_{\pm 2.3}$ $63.0_{\pm 1.8}$ $75.6_{\pm 0.8}$ $73.7_{\pm 1.0}$
Pissa $66.4_{\pm 0.6}$ $80.1_{\pm 0.4}$ $75.6_{\pm 0.9}$ $87.6_{\pm 0.3}$ $63.1_{\pm 0.5}$ $77.6_{\pm 4.0}$ $64.3_{\pm 2.6}$ $75.0_{\pm 1.1}$ $73.7_{\pm 0.9}$
RiemannLoRA $65.9_{\pm 0.7}$ $79.5_{\pm 0.4}$ $76.2_{\pm 0.4}$ $87.4_{\pm 0.3}$ $62.6_{\pm 1.0}$ $78.2_{\pm 1.7}$ $64.3_{\pm 1.7}$ $75.7_{\pm 0.8}$ $73.7_{\pm 0.4}$
RiemannLoRA-LOI $65.8_{\pm 0.2}$ $81.0_{\pm 0.4}$ $\mathbf{76.6}_{\pm 0.4}$ $\mathbf{88.2}_{\pm 0.3}$ $63.7_{\pm 0.8}$ $78.8_{\pm 1.4}$ $\mathbf{64.6}_{\pm 1.7}$ $\mathbf{76.1}_{\pm 0.7}$ $\mathbf{74.3}_{\pm 0.2}$
@@ 5.2 Subject-driven generation
In our experiments, we used Stable Diffusion 2 (Rombach et al., 2022) as the base model. We choose LoRA Hu et al. (2022) as a baseline and train both models with ranks of $4$, $8$ and $16$. We predict the parameterization of the q, k, v, and out.0 matrices in all attention layers. The Dreambooth dataset (Ruiz et al., 2023) was used in all our experiments. LoRA was trained using the Adam optimizer, while our model was trained using our version of the Adam optimizer with the Riemannian optimization (see the Algorithm 3 with simulate_Adam flag enabled).
We used the CLIP model to measure both the text and image similarities of the generated images. Figure 2 shows the training curves of these metrics, where larger dots indicate a higher training step. As can be seen, our method requires fewer steps to learn the concept while preserving the appropriate text similarity. This is also confirmed by the visual comparison presented in Figure 3, for which a training step of $600$ was used to generate images. More quantitative and visual comparisons can be found in the Appendix D and Appendix E.
Figure 2: Comparison of text and image similarities for LoRA and our method at different training steps. Larger dots indicate a higher training step
Figure 3: Visual results for Subject-driven generation on 600 training step
6 Limitations
Since the presented algorithm is a modification of LoRA it inherits the vanilla LoRA limitations, e.g., the problem of underperforming in comparison to full fine-tuning (see (Biderman et al., 2024)), a potentially larger amount of learning parameters in contrast to some other approaches, see, e.g., (Kopiczko et al., 2024). One may also think that our method is constrained by the particular choice of the fixed rank manifold, but generally speaking the approach remains applicable as long as the geometry of the adapter is described by a smooth manifold.
@@ 7 Conclusion
In this work, we propose a new Riemannian framework that addresses the ambiguities inherent in LoRA. In particular, our method is parametrization free and selects the initialization based on the geometric interpretation of Riemannian manifolds. We observe that the proposed method boosts the metric values for different optimization strategies, has small variance and also speeds up convergence in some scenarios. As future work, it would be interesting to explore other optimization strategies and manifolds based on our framework.
@@ Appendix A Inital point search
then one may derive an orthogonal projection formula for any matrix $Z$ (see, e.g. [Boumal, 2023, eq. 7.53])
@@
$\|\nabla\mathcal{L}\|^{2}_{F}=\|\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|^{2}_{F}+\|P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|_{F}^{2},$
@@
Due to the fact that every vector of the tangent space is an element of a set of all matrices with $2r$-bounded rank $\mathcal{M}_{\leq 2r}$, one may use the Eckart-Young-Mirsky theorem (see [Eckart and Young, 1936]) to obtain the following lower bound:
@@
one arrives to 13:
$\begin{split}&\xi_{*}=\begin{bmatrix}A_{*},\dot{A}_{*}\end{bmatrix}\begin{bmatrix}S_{11}&S_{12}\\ S_{21}&S_{22}\end{bmatrix}\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix}\begin{bmatrix}\dot{B_{*}}&B_{*}\end{bmatrix}^{\top},\\ &\Delta W^{(0)}_{*}=\begin{bmatrix}A^{\prime}_{*},\dot{A}^{\prime}_{*}\end{bmatrix}\begin{bmatrix}S_{11}\\ S_{21}\end{bmatrix}\begin{bmatrix}C_{21}&C_{22}\end{bmatrix}\begin{bmatrix}{\dot{B}_{*}^{\prime}}&{{B^{\prime}}_{*}}\end{bmatrix}^{\top}.\end{split}$

### v2

@@ 7 Experiments
In this section, we also employ the term LOI (Locally Optimal Initialization), referring to the proposed initialization scheme described in Section 5. This is done to distinguish between using RiemannLoRA as an optimizer with zero initialization (as in basic LoRA) and RiemannLoRA-LOI, which is the proposed combination of initialization and optimization. We conduct a series of LLM fine-tuning experiments for different tasks. All of the experiments were computed on NVIDIA V100-32Gb GPU and A100-80Gb GPU. We ran all the experiments within $\sim 2000$ GPU hours.
@@ 7.1 Commonsense reasoning fine-tuning
The results were obtained for the benchmark (Clark et al. (2019, BoolQ), Bisk et al. (2020, PIQA), Sap et al. (2019, SIQA), Zellers et al. (2019, hellaswag), Sakaguchi et al. (2021, winogrande),Clark et al. (2018, ARC), Mihaylov et al. (2018, OBQA)) common reasoning. The structure of the dataset is described in Appendix E. In the following experiments we conduct a fine-tuning procedures for multilayer perceptron (MLP) and attention layers of Llama 3 8b model (Dubey et al. (2024)).
@@
The hyperparameter tuning protocol and the selected hyperparameters are provided in F. Table 1 contains the accuracy of the trained model’s responses on the test dataset. Within the LoRA framework, the proposed Riemannian fine-tuning method delivers clear performance gains. The Riemannion optimizer consistently outperforms LoRA, DoRA, and achieves superior metric results compared to the standard Muon optimizer applied to LoRA factors (see Section 3.1). Furthermore, relative to other Riemannian-geometry–aware approaches such as RPrecAdamW (Zhang & Pilanci, 2024), our method also demonstrates better results. Finally, the variance of outcomes for the proposed method is the smallest among all compared approaches.
Table 1: The average accuracy (in %) among 88 tasks of fine-tuned Llama 3-8b using different approaches, tested on Commonsense Reasoning benchmark. LoRA rank is set to 16.
@@
Raw $65.0$ $76.6$ $73.0$ $66.1$ $61.3$ $92.5$ $82.3$ $79.6$ $74.5$
Adam $74.8_{\pm 1.9}$ $89.8_{\pm 0.9}$ $82.6_{\pm 0.6}$ $96.2_{\pm 0.3}$ $87.9_{\pm 1.2}$ $92.4_{\pm 0.7}$ $4.9_{\pm 0.7} 88$$8.5_{\pm 0.4}$ $87.1_{\pm 0.6}$
DoRA $74.8_{\pm 0.8}$ $89.4_{\pm 0.5}$ $82.4_{\pm 0.7}$ $95.9_{\pm 0.1}$ $87.8_{\pm 0.4}$ $90.7_{\pm 1.2}$ $83.8_{\pm 0.7}$ $87.8_{\pm 0.6}$ $86.6_{\pm 0.3}$
Muon $72.9_{\pm 0.0}$ $86.4_{\pm 0.5}$ $80.8_{\pm 0.2}$ $94.1_{\pm 0.2}$ $84.4_{\pm 0.0}$ $84.2_{\pm 0.9}$ $77.3_{\pm 2.5}$ $83.9_{\pm 1.1}$ $83.0_{\pm 0.6}$
DoneRITE $72.2_{\pm 0.3}$ $88.6_{\pm 0.1}$ $82.0_{\pm 0.6}$ $95.1_{\pm 0.1}$ $85.6_{\pm 0.2}$ $87.7_{\pm 1.5}$ $79.3_{\pm 2.6}$ $85.7_{\pm 0.5}$ $84.5_{\pm 0.5}$
RPrecAdamW $\mathbf{75.8}_{\pm 0.4}$ $89.5_{\pm 0.4}$ $82.4_{\pm 0.2}$ $96.1_{\pm 0.2}$ $87.7_{\pm 0.9}$ $90.6_{\pm 1.6}$ $84.1_{\pm 1.1}$ $87.7_{\pm 0.5}$ $86.8_{\pm 0.4}$
Riemannion $75.7_{\pm 0.7}$ $\mathbf{91.2_{\pm 0.2}}$ $\mathbf{83.5_{\pm 0.6}}$ $\mathbf{96.7}_{\pm 0.0}$ $\mathbf{88.6}_{\pm 0.4}$ $\mathbf{93.6}_{\pm 0.3}$ $\mathbf{86.4_{\pm 0.4}}$ $\mathbf{89.3}_{\pm 0.8}$ $\mathbf{88.1}_{\pm 0.2}$
@@ 7.2 Subject-driven generation
In our experiments, we used Stable Diffusion 2 (Rombach et al., 2022) as the base model. We choose LoRA Hu et al. (2022) as a baseline and train both models with ranks of $4$, 88 and $16$.
Figure 1: Comparison of text and image similarities for LoRA and our method with rank 4 at different learning rates on 400 step.
We predict the parameterization of the q, k, v, and out.0 matrices in all attention layers. The Dreambooth dataset (Ruiz et al., 2023) was used in all our experiments. LoRA was trained using the Adam optimizer. In Figure 2, we present a visual comparison of LoRA with different ranks. As can be seen, even for complex concepts such as ’robot toy’, our method requires only 600 steps to learn the concept while preserving appropriate text similarity. We found that for subject-driven generation tasks, the lower the rank, the faster our method converges. To evaluate this, we also calculated metrics for different learning rates on a subset of the Dreambooth dataset. We use CLIP to measure text similarity and DINO to measure image similarity. Figure 1 shows that, even with different learning rates, our method achieves more accurate results in concept preservation. Further details and a visual comparison can be found in the Appendix G.
Figure 2: Visual results for Subject-driven generation on 600 training step.
@@ 8 Conclusion
In this work, we propose a novel fully Riemannian framework that integrates a new muon-based optimization method, locally-optimal initialization, and an efficient implementation. This integrated approach yields a reliable reparametrization-invariant method that outperforms competing approaches on fine-tuning large language models (LLMs) and exhibits additional favorable properties for low-rank approximations in diffusion models. Given the promising empirical results, a natural direction for future research is to investigate the theoretical properties of the proposed method.
9 Reproducibility Statement
The hyperparameter selection procedure is described in Appendix F. The datasets used in the experiments and the corresponding preprocessing steps are detailed in Appendix E. The proof of the stated assumptions is provided in Appendix A. The hyperparameters for Subject-driven generation task are provided in the Appendix G.
@@ Appendix A Inital point search
then one may derive an orthogonal projection formula for any matrix $Z$ (see, e.g. (Boumal, 2023, eq. 7.53))
@@
$\|\nabla\mathcal{L}\|^{2}_{F}$ $=\|\left(\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\right)+P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|^{2}_{F}=$
$=\|\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|^{2}_{F}+\|P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|_{F}^{2}+2\langle\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L},P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\rangle$
$=\|\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|^{2}_{F}+\|P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|_{F}^{2}+2\langle P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\left(\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\right),\nabla\mathcal{L}\rangle$
$=\|\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|^{2}_{F}+\|P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|_{F}^{2}+2\langle\underbrace{P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}}_{=0},\nabla\mathcal{L}\rangle$
$=\|\nabla\mathcal{L}-P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|^{2}_{F}+\|P_{\mathcal{T}_{\Delta W}\mathcal{M}_{r}}\nabla\mathcal{L}\|_{F}^{2}$
@@
Due to the fact that every vector of the tangent space is an element of a set of all matrices with $2r$-bounded rank $\mathcal{M}_{\leq 2r}$, one may use the Eckart-Young-Mirsky theorem (see (Eckart & Young, 1936)) to obtain the following lower bound:
@@
one arrives to (17):
[…]